# Microwave Signatures of Topological Superconductivity in Planar Josephson Junctions

### v1

@@ Abstract
Figure 2: (a), (c) ABS spectra as a function of $\phi$ for JJ from Fig. 1(a). $W_{S}=600\;$nm, $W_{N}=80\;$nm, $L=4\;\mu$m, and $B=0.2\;$T. (b), (d) The corresponding topological charge, $Q_{D}$, as a function of $B$ and $\phi$. (a), (b) For $\mu=0.1\;$meV, the gapped ABS are trivial for all $\phi$. (c) For $\mu=0.1\;$meV, there is a gap closing and reopening. The inset: MBS at $E\approx 0$ (red solid line). (d) A topological transition as $\phi$ is varied at $B=0.2\;$T, corresponding to the gap closing points in (c).
Figure 3: (a), (b) Transition lines $\widetilde{\mathcal{T}}(f,\phi)$, calculated using Eq. (4), corresponding to $hf_{d}=\Delta E=|E_{n}-E_{m}|$ between ABS from Figs. 2(a) and (c), with a Lorentzian broadening parameter $\varepsilon=30\;$mK. In the schematic of pair (c) and single-particle (d) transitions where the ground state is denoted by a blue line, the sum of the arrow lengths in (c) gives $\Delta E$. Resolving (e) pair and (f) single-particle transitions from (a) up to 404040\;40GHz shows no gap closing and opening. In contrast a similarly resolved (g) pair and (h) single-particle transitions from (b) reveal signatures of topological superconductivity. The inset: an enlarged part of the Fig. 2(c) spectra. The transitions in the circles “A” and “B” of the inset lead to the “V”-shaped features “A” and “B” in (g), signaling a gap closing and reopening. Transitions in the circle “C” lead to the “V”-shaped feature “C” in (h).
@@
For our model experimental setup based on fabricated JJs embedded in microwave resonator circuits, we can examine the validity of this approach by a simple estimate, $\delta\phi/2\pi\sim M\,I_{\mathrm{MW}}/\Phi_{0}<M\,I_{c}/\Phi_{0}\approx 0.05$, where we use an experimentaly relevant value of $M\sim 120\;$pH for the mutual inductance and a critical current of $I_{c}\lesssim 1\;\mu$A as an upper bound for the microwave driving current $I_{\mathrm{MW}}$ Dartiailh et al. (2021b); Phan et al. (2023), thus justifying the treatment of the effects of the microwave drive as a perturbation Väyrynen et al. (2015); Wesdorp et al. (2024); Chidambaram et al. (2022). The transitions caused by this perturbation conserve the fermion parity Bretheau et al. (2017); Janvier et al. (2015); Kos et al. (2013); Väyrynen et al. (2015); SM , which reduces the number of the allowed transition lines shown in Figs. 3(a) and (b). We can then distinguish the pair transitions, between a Cooper pair and two ABS, or the single-particle transitions between two ABS, as denoted in Figs. 3(c) and (d). In the BdG Hamiltonian, pair (single-particle) transitions correspond to transitions between the states with opposite (same) energy sign.
@@
By comparing these two types of transitions for the trivial [Figs. 3(e) and (f)] and topological phase [Figs. 3(g) and (h)], we see some striking differences. This is further illustrated in the inset of Fig. 3(g), which is the enlarged version of Fig. 2(c) focusing on low-energy ABS spectra near the topological gap closing and reopening, and the matching features in Fig. 3(g) and (h). The features marked by “A” and “B” in the inset denote the locations of topological gap closing and reopening and the MBS at $E\approx 0$ between these points. The matching “V”-shaped features seen in Fig. 3(g), also marked by “A” and “B,” resolve the pair transition lines around the gap closing and reopening, which therefore mark the microwave signatures of the topological superconductivity. On the other hand, the “V”-shaped single-particle transition lines seen in Fig. 3(h), marked by “C”, correspond to the transitions between the ABS levels in a region far from $E\approx 0$ state (marked by “C” in the inset) and therefore are not related to topological superconductivity. We also note that the features near zero frequency at $\phi\sim\pi/2,3\pi/2$ in Fig. 3(a), (b), (f) and (h) are related to single-particle transitions between ABS near $E\sim\Delta_{0}$, and are not topology related.
@@
We choose a special form of the thermal occupation $\mathcal{F}(E_{m},E_{n})=f_{\mathrm{FD}}(E_{n})\,\left(1-f_{\mathrm{FD}}(E_{m})\right),$ where $f_{\mathrm{FD}}$ is the Fermi-Dirac function, to achieve the effect of suppressing single-particle transitions between the levels away from $E\approx 0$, due to the doubling of the fermionic degrees of freedom in the BdG Hamiltonian. This form of $\mathcal{F}(E_{m},E_{n})$ allows us to simulate the suppression of quasiparticle transitions between energy levels $E_{(}m,n)$ away from 0. The identification of topological signatures is simplified in the resulting transition lines $\widetilde{\mathcal{T}}$ and the shift in the current $\delta I(f,\phi)$. A counterpart plot with a conventional thermal occupation factor is discussed in SM SM_ . In the limit of $T=0\;$K and $\epsilon=0$, our Eq. (6) recovers the same form as in the published literature Blais et al. (2021); Bretheau et al. (2014); Olivares et al. (2014). While Eq. (6) does not include the $\phi$-dependent change of the slope and curvature of the ABS, our key results and trends are preserved by this simplification Park et al. (2020); SM .
For the JJ from Fig. 2(c), with single-particle transitions where $|E|\gtrsim k_{B}T$, we use this approach to plot the transition lines in Figs. 4(a)-(c) and $\delta I(f,\phi)$ of Eq. (6) in Figs. 4(d)-(f). Both the transition lines and $\delta I(f,\phi)$ reveal the signatures of the gap closing and opening, even for a system with hundreds of ABS. While these low-frequency transitions near the gap closing points appear in Figs. 4(b) and (e) for the pair transition, an even more striking feature becomes visible in Figs. 4(c) and (f) for the single-particle transitions. The suppression of the higher-energy single-particle transitions, including the feature “C” in Fig. 3(h), removes the background that competes with the signal from transitions between the energy-split MBS and higher energies, now observed prominently in Fig. 4(c) as a transition line between $\phi\sim 0.8\pi$ and $\phi\sim 1.2\pi$ with half of the slope of the “V” shapes marked by “A” and “C” in Fig. 3(g) Väyrynen et al. (2015); Murthy et al. (2020). This “half-slope” feature is also observed in Figs. 36 in the panels that depict topological transitions. The time-resolved capabilities of microwave spectroscopy could then also provide a detection of the non-Abelian statistics in planar JJs Zhou et al. (2020, 2022), which is supported by experimentally measured parameters, but is very challenging using conventional transport measurements.
@@
Figure 5: (a) ABS spectrum as a function of $B$ showing the lowest 240 levels with $\mu=1$ meV and $\phi=0.75\pi$. The gap closing and reopening, and the appearance of the zero-energy MBS at $B=0.24\;$T match the topological phase transition (see SM SM_ ). (b) Transition lines $\widetilde{\mathcal{T}}$ as a function of $B$. “V” feature indicating that the gap is closing and reopening is visible at $B=0.24\;$T near $f_{d}\sim 0$.
The “V” feature of topological gap closing and reopening is also seen in Fig. 5 from our numerical results for an experimental system where the JJ is phase biased and the Zeeman field is varied. We demonstrate the case where the phase bias is $\phi=0.75\pi$ (at $\phi=\pi$ the transition would be very close to $B=0\;$T). The ABS spectrum in Fig. 5(a) and the topological phase diagram in SM SM_ indicate a transition at $B=0.24\;$T. A corresponding “V” feature in the transition lines is observed in Fig. 5(b). The resolved pair transitions, and the $\delta\,I(f,B)$ plots in SM SM_ also demonstrate the “V” feature and the gap reopening, indicating that topological transitions may be observed in phase-biased microwave experiments.
Figure 6: (a) ABS spectrum as a function of $\phi$ showing the lowest 240 levels with $\mu=5$ meV and $B=0.5\;$T, for $L=4\;\mu$m. (b) Transition lines $\widetilde{\mathcal{T}}$ as a function of $B$. The double “V” feature, indicating the gap closing and reopening, is less sharp but visible around $\phi=0.75\pi$ and $\phi=1.25\pi$ near $f_{d}\sim 0$.
We next examine the influence of a higher chemical potential in a JJ with many ABS levels. The corresponding calculated results are shown in Fig. 6. By comparing the results for the spectra in Figs. 2(c) and 6(a), we see that at a higher $\mu$ will be less defined at a constant $B$ as a function of $\phi$, since a wider range of ABS levels will cross $E=0$. This will result in a reduced topological gap and less protected MBS, and will lead to a less sharp “V”-shaped signature in the transition lines [Fig. 6(b)]. A larger $B$ will shift the gap closing points further apart by shifting the ABS lines in the spectra, making the gap reopening signature more distinct. Similarly, a larger finite $L$ will open the topological gap further, having the same effect of making the signatures more distinct (see SM SM_ ). In our simulations with $\mu=5\;$meV and $L=4\;\mu$m, $\alpha=10\;$meV nm at $B=0.5\;$T, the gap reopening signature remains prominently visible but less sharp in the transition lines.
@@ Acknowledgements.
We thank Prof. Leonid Glazman for valuable discussions. This work is supported by US ONR under awards MURI N000142212764, N000141712793. Computational resources were provided by the UB Center for Computational Research.

### v4

@@ Abstract
Figure 2: (a), (c) ABS spectra as a function of $\phi$ for JJ from Fig. 1(a). $W_{S}=600\;$nm, $W_{N}=80\;$nm, $L=4\;\mu$m, and $B=0.2\;$T. (b), (d) The corresponding topological charge, $Q_{D}$, as a function of $B$ and $\phi$. (a), (b) For $\mu=0.1\;$meV, the gapped ABS are trivial for all $\phi$. (c) For $\mu=1.0\;$meV, there is a gap closing and reopening. The inset: MBS at $E\approx 0$ (red solid line). (d) A topological transition as $\phi$ is varied at $B=0.2\;$T, corresponding to the gap closing points in (c).
Figure 3: (a), (b) Transition lines $\widetilde{\mathcal{T}}(f,\phi)$, calculated using Eq. (4), corresponding to $hf_{d}=\Delta E=|E_{n}-E_{m}|$ between ABS from Figs. 2(a) and (c), with a Lorentzian broadening parameter $\epsilon=30\;$mK. In the schematic of pair (c) and single-particle (d) transitions where the ground state is denoted by a blue line, the sum of the arrow lengths in (c) gives $\Delta E$. Resolving (e) pair and (f) single-particle transitions from (a) up to 404040\;40GHz shows no gap closing and opening. In contrast a similarly resolved (g) pair and (h) single-particle transitions from (b) reveal signatures of topological superconductivity. The inset: an enlarged part of the Fig. 2(c) spectra. The transitions in the circles “A” and “B” of the inset lead to the V-shaped features “A” and “B” in (g), signaling a gap closing and reopening. Transitions in the circle “C” lead to the V-shaped feature “C” in (h).
@@
For our model experimental setup based on fabricated JJs embedded in microwave resonator circuits, we can examine the validity of this approach by a simple estimate, $\delta\phi/2\pi\sim M\,I_{\mathrm{MW}}/\Phi_{0}<M\,I_{c}/\Phi_{0}\approx 0.05$, where we use an experimentaly relevant value of $M\sim 120\;$pH for the mutual inductance and a critical current of $I_{c}\lesssim 1\;\mu$A as an upper bound for the microwave driving current $I_{\mathrm{MW}}$ Dartiailh et al. (2021b); Phan et al. (2023), thus justifying the treatment of the effects of the microwave drive as a perturbation Väyrynen et al. (2015); Wesdorp et al. (2024); Chidambaram et al. (2022). The transitions caused by this perturbation conserve the fermion parity Bretheau et al. (2017); Janvier et al. (2015); Kos et al. (2013); Väyrynen et al. (2015); SM_ , which reduces the number of the allowed transition lines shown in Figs. 3(a) and (b). We can then distinguish the pair transitions, between a Cooper pair and two ABS, or the single-particle transitions between two ABS, as denoted in Figs. 3(c) and (d). In the BdG Hamiltonian, pair (single-particle) transitions correspond to transitions between the states with opposite (same) energy sign.
@@
By comparing these two types of transitions for the trivial [Figs. 3(e) and (f)] and topological phase [Figs. 3(g) and (h)], we see some striking differences. This is further illustrated in the inset of Fig. 3(g), which is the enlarged version of Fig. 2(c) focusing on low-energy ABS spectra near the topological gap closing and reopening, and the matching features in Fig. 3(g) and (h). The features marked by “A” and “B” in the inset denote the locations of topological gap closing and reopening and the MBS at $E\approx 0$ between these points. The matching V-shaped features seen in Fig. 3(g), also marked by “A” and “B,” resolve the pair transition lines around the gap closing and reopening, which therefore mark the microwave signatures of the topological superconductivity. On the other hand, the V-shaped single-particle transition lines seen in Fig. 3(h), marked by “C”, correspond to the transitions between the ABS levels in a region far from $E\approx 0$ state (marked by “C” in the inset) and therefore are not related to topological superconductivity. We also note that the features near zero frequency at $\phi\sim\pi/2,3\pi/2$ in Fig. 3(a), (b), (f) and (h) are related to single-particle transitions between ABS near $E\sim\Delta_{0}$, and are not topology related.
@@
We choose a special form of the thermal occupation $\mathcal{F}(E_{m},E_{n})=f_{\mathrm{FD}}(E_{n})\,\left(1-f_{\mathrm{FD}}(E_{m})\right),$ where $f_{\mathrm{FD}}$ is the Fermi-Dirac function, to achieve the effect of suppressing single-particle transitions between the levels away from $E\approx 0$, due to the doubling of the fermionic degrees of freedom in the BdG Hamiltonian. This form of $\mathcal{F}(E_{m},E_{n})$ allows us to simulate the suppression of quasiparticle transitions between energy levels $E_{m,n}$ away from 0. The identification of topological signatures is simplified in the resulting transition lines $\widetilde{\mathcal{T}}$ and the shift in the current $\delta I(f,\phi)$. A counterpart plot with a conventional thermal occupation factor is discussed in SM SM_ . In the limit of $T=0\;$K and $\epsilon=0$, our Eq. (6) recovers the same form as in the published literature Blais et al. (2021); Bretheau et al. (2014); Olivares et al. (2014). While Eq. (6) does not include the $\phi$-dependent change of the slope and curvature of the ABS, our key results and trends are preserved by this simplification Park et al. (2020); SM_ .
For the JJ from Fig. 2(c), with single-particle transitions where $|E|\gtrsim k_{B}T$, we use this approach to plot the transition lines in Figs. 4(a)-(c) and $\delta I(f,\phi)$ of Eq. (6) in Figs. 4(d)-(f). Both the transition lines and $\delta I(f,\phi)$ reveal the signatures of the gap closing and opening, even for a system with hundreds of ABS. While these low-frequency transitions near the gap closing points appear in Figs. 4(b) and (e) for the pair transition, an even more striking feature becomes visible in Figs. 4(c) and (f) for the single-particle transitions. The suppression of the higher-energy single-particle transitions, including the feature “C” in Fig. 3(h), removes the background that competes with the signal from transitions between the energy-split MBS and higher energies, now observed prominently in Fig. 4(c) as a transition line between $\phi\sim 0.8\pi$ and $\phi\sim 1.2\pi$ with half of the slope of the V shapes marked by “A” and “C” in Fig. 3(g) Väyrynen et al. (2015); Murthy et al. (2020). This “half-slope” feature is also observed in Figs. 36 in the panels that depict topological transitions. The time-resolved capabilities of microwave spectroscopy could then also provide a detection of the non-Abelian statistics in planar JJs Zhou et al. (2020, 2022), which is supported by experimentally measured parameters, but is very challenging using conventional transport measurements.
@@
Figure 5: (a) ABS spectrum as a function of $B$ showing the lowest 240 levels with $\mu=1$ meV and $\phi=0.75\pi$. The gap closing and reopening, and the appearance of the zero-energy MBS at $B=0.24\;$T match the topological phase transition (see SM SM_ ). (b) Transition lines $\widetilde{\mathcal{T}}$ as a function of $B$. V feature indicating that the gap is closing and reopening is visible at $B=0.24\;$T near $f_{d}\sim 0$.
The V feature of topological gap closing and reopening is also seen in Fig. 5 from our numerical results for an experimental system where the JJ is phase biased and the Zeeman field is varied. We demonstrate the case where the phase bias is $\phi=0.75\pi$ (at $\phi=\pi$ the transition would be very close to $B=0\;$T). The ABS spectrum in Fig. 5(a) and the topological phase diagram in SM SM_ indicate a transition at $B=0.24\;$T. A corresponding V feature in the transition lines is observed in Fig. 5(b). The resolved pair transitions, and the $\delta\,I(f,B)$ plots in SM SM_ also demonstrate the V feature and the gap reopening, indicating that topological transitions may be observed in phase-biased microwave experiments.
Figure 6: (a) ABS spectrum as a function of $\phi$ showing the lowest 240 levels with $\mu=5$ meV and $B=0.5\;$T, for $L=4\;\mu$m. (b) Transition lines $\widetilde{\mathcal{T}}$ as a function of $B$. The double V feature, indicating the gap closing and reopening, is less sharp but visible around $\phi=0.75\pi$ and $\phi=1.25\pi$ near $f_{d}\sim 0$.
We next examine the influence of a higher chemical potential in a JJ with many ABS levels. The corresponding calculated results are shown in Fig. 6. By comparing the results for the spectra in Figs. 2(c) and 6(a), we see that at a higher $\mu$ will be less defined at a constant $B$ as a function of $\phi$, since a wider range of ABS levels will cross $E=0$. This will result in a reduced topological gap and less protected MBS, and will lead to a less sharp V-shaped signature in the transition lines [Fig. 6(b)]. A larger $B$ will shift the gap closing points further apart by shifting the ABS lines in the spectra, making the gap reopening signature more distinct. Similarly, a larger finite $L$ will open the topological gap further, having the same effect of making the signatures more distinct (see SM SM_ ). In our simulations with $\mu=5\;$meV and $L=4\;\mu$m, $\alpha=10\;$meV nm at $B=0.5\;$T, the gap reopening signature remains prominently visible but less sharp in the transition lines.
@@
Note added–Recently, we were made aware of a publication Trif et al. (2018) titled “Current susceptibility as a probe of Majorana bound states in nanowire-based Josephson junctions,” which is relevant to our work.
@@ Acknowledgements.
We thank Prof. Leonid Glazman for valuable discussions. This work is supported by the U.S. ONR MURI Award No. N000142212764 and U.S. ONR Award No. N000141712793. Computational resources were provided by the UB Center for Computational Research.